# Convolutional autoencoders for the reconstruction of three-dimensional interfacial multiphase flows

### v1

@@ Abstract
In this work, we perform a comprehensive investigation of autoencoders for reduced-order modeling of three-dimensional multiphase flows. Focusing on the accuracy of reconstructing multiphase flow volume/mass fractions with a standard convolutional architecture, we examine the advantages and disadvantages of different interface representation choices (diffuse, sharp, level set). We use a combination of synthetic data with non-trivial interface topologies and high-resolution simulation data of multiphase homogeneous isotropic turbulence for training and validation. This study clarifies the best practices for reducing the dimensionality of multiphase flows via autoencoders. Consequently, this paves the path for uncoupling the training of autoencoders for accurate reconstruction and the training of temporal or input/output models such as neural operators (e.g., FNOs, DeepONets) and neural ODEs on the lower-dimensional latent space given by the autoencoders. As such, the implications of this study are significant and of interest to the multiphase flow community and beyond.
@@
autoencoder , multiphase flow , reduced order model , deep learning , interface capturing
@@ 1 Introduction
Fluid flows are considered high-dimensional dynamical systems, where real-time prediction, control, and optimization are prohibitively expensive without some form of reduced-order modeling. A low-dimensional representation of the state of a fluid system is required for a reduced-order model (ROM). While there has been substantial progress in developing ROMs for single-phase flows [1, 2, 3, 4], interfacial multiphase flows present a unique challenge. Multiphase flows consist of two or more immiscible fluid phases separated by a deformable interface. In this case, the state of the dynamical system strongly depends on the instantaneous location of the different phases, due to the discontinuity in density, viscosity, pressure, and other fields across the deformable phase interface. As such, it is crucial for the ROM to represent the shape of three-dimensional interfaces in multiphase flows accurately. In this work, we study the performance of convolutional autoencoders (AE) for this task. We reveal the impact of the interface representation type (diffuse, sharp, level-set function), shape complexity, training set set size, and compression ratio on the performance of the AE. This work constitutes a robust building block for developing ROMs to predict temporal dynamics in multiphase flows. Finally, considering that low-dimensional representations of complex three-dimensional shapes are highly sought after in many applications, including computer graphics, computer vision, and robotics, our findings may have implications beyond multiphase flow modeling [5, 6, 7].
@@
While substantial progress has been made in simulating (full-order modeling) interfacial multiphase flows using various approaches in the past few decades [8, 9], these simulations are too expensive to allow for real-time predictions, control, or optimization. As such, ROMs for such multiphase flows are sought after. Such models are scarce in the literature [10, 11, 12, 13, 14], especially when one seeks ROMs that aim to capture the interfaces and their evolution in such flows. Here, we focus on the low-dimensional representation of the interfaces in multiphase flows via convolutional autoencoders (AEs). An AE is a neural network consisting of an encoder that maps the input onto a latent representation and a decoder that reconstructs the input from this representation. The encoder performs a nonlinear transformation, often reducing dimensionality, while the decoder attempts to approximate the original input as closely as possible. AEs have been successful in low-dimensional representation of single-phase flows [15, 16, 17, 18], but their application to the reconstruction of multiphase flow states remains limited [10, 13]. There are multiple ways to implicitly represent interfaces in multiphase flows via field variables. These include diffuse interface representations like phase field variables, sharp interface representations like volume fractions in Volume-of-Fluid models, and level-set functions (signed-distance functions), which are inherently available in level-set-based multiphase flow models. Since one can transform these representations into one another (e.g., as a pre-processing step), we explore how the interface representation choice influences the performance of the AE. For training, we use a series of synthetic data sets with increasing interface complexity, as well as a data set consisting of snapshots from high-fidelity simulations of drops in decaying homogeneous isotropic turbulence (HIT) [14]. This also allows us to study the effect of the complexity of the interfacial structures on the reconstruction accuracy of the AEs. All in all, accurately representing flow states in a low-dimensional space is a prerequisite for most ROMs. This work elucidates the best practices and limitations of applying AEs for this purpose in multiphase flows.
@@ 2.1 Problem formulation
We seek to train convolutional autoencoders (AEs) for the reconstruction of interfacial structures in two-phase systems. Interfacial structures can be captured with various field variables. Figure 1 shows three different choices for full-order representation of the same interfacial structure, where panel (a) displays the signed distance function (SDF), also known as the level-set function, typically denoted with $s$. Panel (b) displays a diffuse interface representation using a so-called phase field variable given by
@@ 2.2.1 Interfacial flow simulation dataset
To incorporate a realistic flow, we use snapshots from simulations of drops in decaying homogeneous isotropic turbulence (HIT) that we previously used for training a machine learning (ML) model for drop breakup prediction [14]. The simulations were performed on a $256^{3}$ grid using a phase field (diffuse interface) approach [19, 20]. Patches of size $64^{3}$ were extracted from these snapshots, by randomly choosing 64 patch locations per volume and discarding empty patches. Knowing the interface thickness ($\epsilon_{\text{sim}}$) used in the simulations, we compute the SDF ($s$ in Section 2), from which we construct input fields with various diffuse interface thicknesses, given by
@@
with the limit of $\epsilon\to 0$ giving us the sharp interface representation.
@@ 2.2.2 Synthetic dataset
$N_{D}$ is then drawn from a uniform distribution such that the total volume fraction is uniformly distributed in $[0,1]$. All samples are $64^{3}$.
@@ 2.3 Neural network architecture
Throughout this work, we adopt a standard convolutional architecture; our objective is to compare interface representations rather than architecture optimization. The choice of a fully convolutional architecture [21] is motivated by our application to (potentially large) 3D volumes, as it enables a patch-based approach due to the translation equivariance of convolutional layers. Our network architecture follows the standard encoder-decoder structure described in Equation 1 and is based on the residual block illustrated in Figure 3. Each layer in the encoder (decoder) consists of two such blocks in series, followed by a downsampling (upsampling) operation. The choice of a ResNet-type architecture [22] along with weight-standardized convolution layers is advantageous in the small batch size regime [23], which is necessarily the case in 3D due to the large feature maps which must be stored.
Given an input of shape $1\times H\times W\times D$, we obtain a latent representation of shape $Z\times\frac{H}{2^{N}}\times\frac{W}{2^{N}}\times\frac{D}{2^{N}}$, where $N$ is the number of downsampling layers in the encoder. Unless otherwise specified, we use $N=4$ and $Z=4$, resulting in a compression ratio of $1024$ and approximately $5.4\times 10^{6}$ trainable parameters.
@@ 2.4 Training
All models were trained using the Adam optimization algorithm [24] and mean absolute error loss, with a batch size of 4 and a learning rate of $10^{-5}$ (see hyper-parameter study in Section 3.3). The number of training epochs was held constant at 100 for the synthetic datasets and 15 for the HIT simulation dataset. This was confirmed qualitatively to provide sufficient convergence. Each training run utilised 4 Nvidia V100 (16GB) GPUs in a data-parallel approach.
@@ 3.1 Evaluation metrics
Each case is evaluated using a held-out test set, according to the splits described in Section 2.2. Rather than standard $L^{p}$ error norms, we adopt two metrics from the field of image segmentation. The evaluation of a (semantic) image segmentation is highly similar to the evaluation of an interfacial geometry. In both cases, the prediction consists of an arbitrary region represented by a binary mask. We transform all predictions to a sharp interface representation before computing metrics. The following two metrics offer a concise and understandable summary of the performance of this prediction.
First, the Dice coefficient [25] provides a value in $[0,1]$ describing the overall volumetric agreement, where a value of 1 is the best possible score:
@@
Second, the Hausdorff distance [26] is a worst-case measure, analogous to the $L^{\infty}$ norm. It is the maximum distance between the predicted and the true interface:
@@ 3.2 Training process uncertainty
Table 1: Measurement of training uncertainty. Mean, standard deviation and 95% confidence interval (CI) for Dice and Hausdorff distance, based on repeats with five random seeds. 50% of HIT simulation dataset used.
@@ 3.3 Hyper-parameter study
A limited hyper-parameter grid search was run using the HIT simulation dataset. The search space consisted of learning rate, loss function, weight decay, and activation function and is summarized in Table 2. The dice scores on the hyper-parameter validation split for three interface types (sharp, tanh $\frac{1}{32}$, SDF) are summarized by a parallel coordinate plot in Figure 4.
@@
The optimal weight decay and activation functions are unambiguously $10^{-6}$ and SiLU. The optimal learning rate is $10^{-5}$, but the sharp interface representation is also tolerant of a $10^{-4}$ learning rate. The choice of loss function appears to be interface-type dependent, as shown in Figure 5. For sharp representations, both MSE and $L_{1}$ are within the uncertainties estimated in the previous section. However for the more diffuse tanh interface representations and the SDF, using the $L_{1}$ loss offers a distinct advantage. This result may be understood in terms of MSE loss providing a smaller signal close to the interface for smoother representations. Further details on the hyperparameter search are provided in A.
@@
Figure 4: Parallel coordinate plot showing hyper-parameter grid search results for three different interface representations.
@@
Figure 6: Absolute performance of each interface representation on (a) HIT simulation dataset and (b) synthetic datasets. Recall that for Dice, higher is better, while for Hausdorff, lower is better.
(a) HIT simulation dataset
(b) Synthetic datasets
Figure 7: Relative performance of each interface representation on (a) HIT simulation dataset and (b) synthetic datasets. Recall that for Dice, higher is better, while for Hausdorff, lower is better.
@@
Figures 6 and 7 show the absolute and relative reconstruction performance for each dataset using various interface representations. The Dice coefficient is generally lowest for the SDF inputs and steadily improves as the interface becomes sharper, plateauing at a dataset-dependent interface thickness $\epsilon$.
@@
While the Dice metric favors sharper interfaces, a different trend emerges when evaluating the Hausdorff distance. For more complex fields containing small drops, the sharpest representations no longer yield the best performance. Instead, a moderately diffuse interface representation with $\epsilon\approx\frac{1}{32}$ tends to give the lowest Hausdorff error. A representative example shown in Figure 8 illustrates this effect: sharper interfaces enable more accurate reconstruction of large-scale features, but at the expense of small-scale ones. We attribute this trade-off to the spectral bias of deep neural networks [27], which makes high-frequency features (such as sharp interfaces and fine structures) more difficult to learn during training.
@@
The droplet size distribution is an important quantity of interest in many interfacial multiphase flows. In Figure 9, we compare the predicted distribution of droplet size in the reconstructed samples on the synthetic dataset. In particular, when we focus on the most complex dataset ($\mu=1$), we observe that the best match to the ground truth was given by a tanh profile with an intermediate interface thickness. This observation further supports the conclusion that moderately diffuse interfaces offer the best balance between resolving large and small features.
@@
The underperformance of SDFs was unexpected and highlights an important distinction between interface representations. We hypothesize that this is because reconstruction errors in the SDF representation are distributed throughout the domain, rather than being localized near the interface, and thus do not penalize the model as effectively as errors in sharp or diffuse representations. Overall, our findings suggest a trade-off between sharp and diffuse representations: while sharper interfaces improve the reconstruction of large-scale features, moderately diffuse representations better capture small-scale structures in complex fields. Despite this trade-off, both sharp and diffuse representations consistently outperform SDFs. We repeated these experiments at compression ratios ranging from 100 to 8000 (not shown for brevity), and observed that these trends persist across compression levels.
@@
3.5 Performance vs. dataset size
For each interface representation and dataset, we investigate the effect of progressively reducing the training set size. As shown in Figure 10, the model performance on less complex datasets continues to improve with more data, converging toward perfect reconstruction at a rate of $\mathcal{O}(N^{-\frac{1}{2}})$. In contrast, for the most complex datasets ($\mu=1$ and $\mu=2$), the model architecture, rather than the dataset size, appears to be the limiting factor.
(a) HIT simulation dataset
(b) Synthetic datasets
Figure 10: Performance versus training set size for each dataset.

### v2

@@ Abstract
We present a systematic investigation of convolutional autoencoders for the reduced-order representation of three-dimensional interfacial multiphase flows. Focusing on the reconstruction of phase indicators, we examine how the choice of interface representation, including sharp, diffuse, and level-set formulations, impacts reconstruction accuracy across a range of interface complexities. Training and validation are performed using both synthetic datasets with controlled geometric complexity and high-fidelity simulations of multiphase homogeneous isotropic turbulence. We show that the interface representation plays a critical role in autoencoder performance. Excessively sharp interfaces lead to the loss of small-scale features, while overly diffuse interfaces degrade overall accuracy. Across all datasets and metrics considered, a moderately diffuse interface provides the best balance between preserving fine-scale structures and achieving accurate reconstructions. These findings elucidate key limitations and best practices for dimensionality reduction of multiphase flows using autoencoders. By clarifying how interface representations interact with the inductive biases of convolutional neural networks, this work lays the foundation for decoupling the training of autoencoders for accurate state compression from the training of surrogate models for temporal forecasting or input-output prediction in latent space.
@@
reduced order model , surrogate model , machine learning , autoencoder , multiphase flow , interface capturing
@@ 1 Introduction
Fluid flows are considered high-dimensional dynamical systems, where real-time prediction, control, and optimization are prohibitively expensive without some form of reduced-order modeling. A low-dimensional representation of the state of a fluid system is required for a reduced-order model (ROM). While there has been substantial progress in developing ROMs for single-phase flows [rowley2017model, brunton2020machine, vinuesa2022enhancing, rozza2022advanced], interfacial multiphase flows present a unique challenge. Multiphase flows consist of two or more immiscible fluid phases separated by a deformable interface. In this case, the state of the dynamical system strongly depends on the instantaneous location of the different phases, due to the discontinuity in density, viscosity, pressure, and other fields across the deformable phase interface. As such, it is crucial for the ROM to represent the shape of three-dimensional interfaces in multiphase flows accurately. In this work, we study the performance of convolutional autoencoders (AE) for this task. We reveal the impact of the interface representation type (diffuse, sharp, level-set function) and shape complexity on the performance of the AE. This work constitutes a robust building block for developing ROMs to predict temporal dynamics and input-output mappings in multiphase flows. Finally, considering that low-dimensional representations of complex three-dimensional shapes are highly sought after in many applications, including computer graphics, computer vision, and robotics, our findings may have implications beyond multiphase flow modeling [bagautdinov2018modeling, dai2017shape, groueix2018papier].
@@
While substantial progress has been made in simulating (full-order modeling) interfacial multiphase flows using various approaches in the past few decades [tryggvason2011direct, mirjalili2017interface], these simulations are too expensive to allow for real-time predictions, control, or optimization. As such, ROMs for such multiphase flows are sought after. Such models are scarce in the literature [wiewel2019latent, kani2019reduced, haas2020bubcnn, zhu2022review, cundy2024physics], especially when one seeks ROMs that aim to capture the interfaces and their evolution in such flows. Here, we focus on the low-dimensional representation of the interfaces in multiphase flows via convolutional AEs. An AE is a neural network consisting of an encoder that maps the input onto a latent representation and a decoder that reconstructs the input from this representation. The encoder performs a nonlinear transformation, often reducing dimensionality, while the decoder attempts to approximate the original input as closely as possible. AEs have been successful in low-dimensional representation of single-phase flows [solera2024beta, brunton2022data, kneer2021symmetry, csala2022comparing], but their application to the reconstruction of multiphase flow states remains limited [wiewel2019latent, zhu2022review]. There are multiple ways to implicitly represent interfaces in multiphase flows via field variables. These include diffuse interface representations like phase field variables, sharp interface representations like volume fractions in Volume-of-Fluid models, and level-set functions (signed-distance functions), which are inherently available in level-set-based multiphase flow models. Since one can transform these representations into one another (e.g., as a pre-processing step), we explore how the interface representation choice influences the performance of the AE. For training, we use a series of synthetic data sets with increasing shape complexity, as well as a data set consisting of snapshots from high-fidelity simulations of drops in decaying homogeneous isotropic turbulence (HIT) [cundy2024physics]. This also allows us to study the effect of the complexity of the interfacial structures on the reconstruction accuracy of the AEs.
While autoencoders have been increasingly adopted in fluid mechanics, existing studies have largely focused on smooth bulk fields such as pressure or velocity [wiewel2019latent], where it has been observed that small-scale features are systematically under-resolved. In interfacial multiphase flows, however, the interface geometry itself constitutes the dominant source of high-frequency content and topological complexity. The main novelty of the present work is a quantitative and systematic assessment of how the choice of interface representation influences the ability of convolutional autoencoders to represent interfacial geometry. By comparing sharp, signed-distance, and diffuse interface representations across datasets with controlled geometric complexity, we identify spectral bias as the key mechanism limiting reconstruction accuracy and demonstrate that moderately diffuse interfaces alleviate this limitation by regularizing high-wavenumber content without compromising geometric fidelity. To the best of our knowledge, this trade-off between interface representation, spectral bias, and reconstruction accuracy has not previously been established for multiphase flows. All in all, accurately representing flow states in a low-dimensional space is a prerequisite for most ROMs. This work elucidates the best practices and limitations of applying AEs for this purpose in multiphase flows.
@@ 2.1 Problem formulation
We seek to train convolutional AEs for the reconstruction of interfacial structures in two-phase systems. Interfacial structures can be captured with various field variables. Figure 1 shows three different choices for full-order representation of the same interfacial structure, where panel (a) displays the signed distance function (SDF), also known as the level-set function, typically denoted with $s$. Panel (b) displays a diffuse interface representation using a so-called phase field variable given by
@@ 2.2.1 Interfacial flow simulation dataset
To incorporate a realistic flow, we use snapshots from simulations of drops in decaying homogeneous isotropic turbulence (HIT) that we previously used for training a machine learning (ML) model for drop breakup prediction [cundy2024physics].
To generate the HIT dataset in  [cundy2024physics], direct numerical simulations (DNS) of drops in decaying homogeneous isotropic turbulence are performed using our in-house two-phase solver based on the conservative Allen–Cahn phase field method [mirjalili2020conservative, mirjalili2021consistent]. The governing parameters are $\rho_{1}=\rho_{2}=5$, $\mu_{1}=\mu_{2}=0.005$, and $\sigma=0.025$ in SI units, corresponding to unity density and viscosity ratios. The resulting flows are characterized by a small Ohnesorge number and Weber numbers in the range $\mathcal{O}(0.1)$$\mathcal{O}(100)$, which produce a wide variety of drop deformation and breakup behaviors relevant to emulsions. All simulations are conducted in a periodic domain of size $1\times 1\times 1$ on a uniform staggered grid of $256^{3}$ cells. The velocity field is initialized by running a single-phase forced HIT simulation for $\Delta t=2$ using the linear forcing method of Rosales and Meneveau (2005), with the forcing coefficient $B$ sampled from $U(1,6.5)$. This produces flows with $Re_{\lambda}\sim\mathcal{O}(10)$ while ensuring that the Kolmogorov scale is resolved. After the forcing is removed, droplets with randomized geometry and position are inserted, and the two-phase system is evolved under decaying turbulence until $t_{\mathrm{final}}=2$. Because the density and viscosity ratios are unity, drop insertion perturbs the momentum field only through surface tension forces, enabling stable two-phase DNS. This workflow allows us to generate on the order of $10^{4}$ independent simulations. For each simulation, $5$ snapshots are saved.
Patches of size $64^{3}$ are extracted from these snapshots, by randomly choosing 64 patch locations per volume and discarding empty patches. Knowing the interface thickness ($\epsilon_{\text{sim}}$) used in the simulations, we compute the SDF ($s$ in Section 2, by inverting Eq. (1)), from which we construct input fields with various diffuse interface thicknesses, given by
@@
with the limit of $\epsilon\to 0$ giving us the sharp interface representation. A total of 25,000 patches are extracted from this dataset.
@@ 2.2.2 Synthetic dataset
The number of droplets $N_{D}$ is drawn from a uniform distribution, $N_{D}\sim U(1,N_{\max})$, where $N_{\max}$ is determined using a reference droplet volume based on the expected radius $\bar{r}=\exp(\mu+0.25)$. Specifically, defining a reference volume $V=\frac{4\pi}{3}\bar{r}^{3}$, we choose $N_{\max}$ such that $N_{\max}V=0.25$. As a result, the realized volume fraction varies stochastically between samples and is not enforced exactly, but spans a broad range up to approximately $0.25$. All samples are $64^{3}$. A total of 10,000 volume patches are generated for each value of $\mu$ in the synthetic datasets.
@@ 2.3 Neural network architecture
Throughout this work, we adopt a standard convolutional architecture [long2015fully] [^2] Our architecture is available at https://github.com/murraycutforth/conv-ae-3d; our objective is to compare interface representations rather than architecture optimization. Our network architecture follows the standard encoder-decoder structure described in Equation 1 and is based on the residual block illustrated in Figure 3. Each layer in the encoder (decoder) consists of two such blocks in series, followed by a downsampling (upsampling) operation. Downsampling in the encoder is achieved via a 3D convolution with a stride of 2, and upsampling in the decoder is performed with a 3D transposed convolution with a stride of 2. All other convolutions use a stride of 1 with padding to preserve spatial dimensions. The choice of a ResNet-type architecture [he2016deep] along with weight-standardized convolution layers is advantageous in the small batch size regime [qiao2019micro], which is necessarily the case in 3D due to the large feature maps which must be stored. A specifies the output shape of each block throughout the network. Given an input of shape $1\times H\times W\times D$, we obtain a latent representation of shape $Z\times\frac{H}{2^{N}}\times\frac{W}{2^{N}}\times\frac{D}{2^{N}}$, where $N$ is the number of downsampling layers in the encoder. Unless otherwise specified, we use $N=4$ and $Z=4$, resulting in a compression ratio of $1024$ and approximately $5.4\times 10^{6}$ trainable parameters.
@@ 2.4 Training
All models were trained using the Adam optimization algorithm [kingma2014adam] and mean absolute error loss, with a batch size of 4 and a learning rate of $10^{-5}$ (see hyper-parameter study in Section 3.3). The number of training epochs was held constant at 100 for the synthetic datasets and 15 for the HIT simulation dataset. This was confirmed qualitatively to provide sufficient convergence. Some exaples of training loss curves are provided in B. Each training run utilised 4 Nvidia V100 (16GB) GPUs for 12 hours in a data-parallel approach. Note that inference cost is much lower and is performed locally, taking 10s per volume on an Apple Macbook Pro M3.
@@ 3.1 Evaluation metrics
[…]
First, the Dice coefficient [dice1945measures] provides a value in $[0,1]$ describing the overall volumetric agreement, where a value of 1 is the best possible score:
@@
Second, the Hausdorff distance [huttenlocher1993comparing] is a worst-case measure, analogous to the $L^{\infty}$ norm. It is the maximum distance between the predicted and the true interface:
@@
Finally, we also measure the relative volume error of the dispersed phase:
$\mathrm{RVE}_{\mathrm{abs}}(X,Y)=\frac{\big||X|-|Y|\big|}{|Y|}.$ (8)
@@ 3.2 Training process uncertainty
Table 1: Measurement of training uncertainty. Mean, standard deviation and 95% confidence interval (CI) for Dice and Hausdorff distance, based on repeats with five random seeds. 50% of HIT simulation dataset used. 1/32{1}/{32} denotes the epsilon (interface width) used with the tanh representation.
@@ 3.3 Hyper-parameter study
A limited hyper-parameter grid search was run using the HIT simulation dataset. The search space consisted of learning rate, loss function, weight decay, and activation function and is summarized in Table 2. The dice scores on the hyper-parameter validation split for three interface types (sharp, tanh $1/32$, SDF) are summarized by a parallel coordinate plot in Figure 4.
@@
The optimal weight decay and activation functions are unambiguously $10^{-6}$ and SiLU. The optimal learning rate is $10^{-5}$, but the sharp interface representation is also tolerant of a $10^{-4}$ learning rate. The choice of loss function appears to be interface-type dependent, as shown in Figure 5. For sharp representations, both MSE and $L_{1}$ are within the uncertainties estimated in the previous section. However for the more diffuse tanh interface representations and the SDF, using the $L_{1}$ loss offers a distinct advantage. This result may be understood in terms of MSE loss providing a smaller signal close to the interface for smoother representations. Further details on the hyperparameter search are provided in C.
@@
Figure 4: Parallel coordinate plot showing hyper-parameter grid search results for three different interface representations. Each line corresponds to a single hyper-parameter set, colored by validation set performance. Line thickness is also proportional to performance in order to highlight the best hyperparameter sets.
@@
Figure 6: Performance of each interface representation on (a) HIT simulation dataset and (b) synthetic datasets. Recall that for Dice, higher is better, while for Hausdorff and relative volume fraction error, lower is better. It was computationally infeasible to repeat all experiments to measure training uncertainty, but note that a subset of these runs were repeated as shown in Table 1, and these suggest that these differences are significantly greater than the training uncertainty.
@@
Figure 6 shows the reconstruction performance for each test dataset using various interface representations. The Dice coefficient is generally lowest for the SDF inputs and steadily improves as the interface becomes sharper, plateauing at a dataset-dependent interface thickness $\epsilon$.
@@
While the Dice metric favors sharper interfaces, a different trend emerges when evaluating the Hausdorff distance. For more complex fields containing small drops, the sharpest representations no longer yield the best performance. Instead, a moderately diffuse interface representation tends to give the lowest Hausdorff error. A representative example shown in Figure 7 illustrates this effect: sharper interfaces enable more accurate reconstruction of large-scale features, but at the expense of small-scale ones. We attribute this trade-off to the spectral bias of deep neural networks [rahaman2019spectral], which makes high-frequency features (such as sharp interfaces and fine structures) more difficult to learn during training. This interpretation is quantitatively supported by a spectral analysis of the reconstructed fields and reconstruction errors performed in D, revealing a systematic attenuation of high-wavenumber content in the autoencoder outputs, particularly for sharp interface representations.
Volume conservation errors for the dispersed phase are also shown in Figure 6. For most datasets, the SDF and the smoothest diffuse interface representation exhibit the largest volume conservation errors. The magnitude of the conservation errors are dataset dependent, typically increasing with dataset complexity. Nonetheless, consistent with the other error metrics, a moderately diffuse interface representation yields comparatively small errors across all datasets.
@@
The droplet size distribution is an important quantity of interest in many interfacial multiphase flows. We compute droplet size distributions by first converting the reconstructed fields to a sharp interface representation and then applying connected-component labeling to identify individual droplets. The volume of each droplet is computed by integrating the indicator function over each connected component. In Figure 8, we compare the predicted distribution of droplet size in the reconstructed samples on the synthetic dataset. In particular, when we focus on the most complex dataset ($\mu=1$), we observe that the best match to the ground truth was given by a tanh profile with an intermediate interface thickness. This observation further supports the conclusion that moderately diffuse interfaces offer the best balance between resolving large and small features.
@@
The underperformance of SDFs was unexpected and highlights an important distinction between interface representations. We hypothesize that this is because reconstruction errors in the SDF representation are distributed throughout the domain, rather than being localized near the interface, and thus do not penalize the model as effectively as errors in sharp or diffuse representations. Overall, our findings suggest a trade-off between sharp and diffuse representations: while sharper interfaces improve the reconstruction of large-scale features, moderately diffuse representations better capture small-scale structures in complex fields. Despite this trade-off, both sharp and diffuse representations consistently outperform SDFs. Finally, to assess whether nonlinear representations are essential for this task, we evaluate a linear autoencoder baseline with the same compression ratio (E). Across all datasets and interface representations, the linear model exhibits substantially larger reconstruction errors than the nonlinear convolutional autoencoder, confirming that nonlinear feature extraction is required to accurately represent complex interfacial geometries.